a researcher wishes to estimate the proportion of adults who have high spped internet access . what size sample should be obtained if she wishes the estimate to be within 0.05 with 90% confidence if she uses a previous estimate of 0.46
Answer #1

Solution:

Sample size formula is given as below:

n = p*q*(Z/E)^2

We are given

p = 0.46

q = 1 – p = 1 – 0.46 = 0.54

Confidence level = 90%

Critical Z value = 1.6449

(by using z-table)

E = 0.05

n = 0.46*0.54*(1.6449/0.05)^2

n = 268.838

Required sample size = 269